# What Should Embeddings Embed? Autoregressive Models Represent Latent Generating Distributions

### v1

@@ 3.5 Probing embeddings to recover predictive sufficient statistics
The three cases identified above specify information that should be encoded in the embeddings formed in neural networks. This sets up the second element of our approach, which is building probes to check whether this information is actually encoded in the representations of specific networks. We focus on transformers (Vaswani et al., 2017), since they are widely used as language models. We denote the target to be decoded for each sequence $x_{1:n}$ by a vector $t$. For example, in the discrete hypothesis space case the target $t$ is $p(h|x_{1:n})$, which is a vector on the simplex with vertices corresponding to different values of $h$. Exact choices of $t_{n}$ in each dataset are given in the next section. Given a trained transformer, we decode $t$ by by training a second model (a probe) to predict the target $t$ from the embedding $\phi_{n}$ of the corresponding document $x_{1:n}$. The probe $g$ maps from the document embedding to the target. To ensure that the relevant statistical information is contained in the LM, not in the probe, we keep the probe simple by defining it as a linear layer with softmax activations: $g(\phi_{n})=\text{Softmax}(\text{Linear}(\phi_{n}))$.
4 Empirical analysis of embeddings
We have identified three cases where a predictive sufficient statistic is expected to be contained in an autoregressive model’s embeddings. In this section, we conduct probing on transformers trained on different datasets to empirically verify this hypothesis. We also conduct ablation studies to further test this idea. In all experiments, we use a linear probe unless otherwise mentioned, three-layer transformer decoder, and the Adam optimizer (Kingma and Ba, 2014) (more implementational details in Appendix A.3).
4.1 Exchangeable models
@@ 4.1.1 Generative model
We evaluate our approach on three exchangeable models: the Gaussian-Gamma, Beta-Bernoulli, and Gamma-Exponential models. Due to space limitations we only describe the Gaussian-Gamma model in detail, but the full generative process underlying the other two models is given in Appendix A.1.
The Gaussian-Gamma model generates from a Gaussian distribution with two unknown parameters. For $M$ sequences $\{x^{(k)}\}_{1\leq k\leq M}$ with $N$ tokens each, each sequence is i.i.d. generated by a mean and precision parameter sampled from a prior. Across the $M$ sequences, $M$ mean and precision parameters are sampled to generate these sequences. For sequence ${x}^{(k)}$, the generative process is,
@@
where $\alpha_{0},\beta_{0},\mu_{0},\lambda_{0}$ are fixed hyperparameters. The predictive distribution for the next token in sequence ${x}$ is,
@@
The optimal Bayesian agent uses the same prior distributions as the Gamma and Normal that generate $\mu$ and $\tau$. It will analytically infer the ground truth posterior $p(\mu,\tau|x_{1:n})$ for any stream of data it sees, and use this posterior for predicting next tokens. A Bayesian agent can also use other suitable priors and converge to the optimal posterior. To be consistent with other exchangeable conjugate models, we denote $\theta=(\mu,\tau)$ to indicate latent variables whose posterior distribution is predictive sufficient.
@@
Figure 2: Probe recovery of transformer-learned sufficient statistic (blue) and ground truth sufficient statistic (red), across 1000 test datapoints. The first row shows parameters probed in the non-OOD case (from left to right: Gaussian mean $\mu$, Gaussian precision $\tau$, Bernoulli mean, and Exponential mean). The second row shows the corresponding information in the OOD case.
@@ 4.1.2 Probing experiments
Experimental setup
We hypothesize that a transformer trained on this dataset should come to represent the sufficient statistics of the corresponding distribution. We use three probing setups.
Sufficient statistic. As suggested by the Proposition in Section 3.2, we probe the sufficient statistic for $\theta$. The probe uses the last token embedding of the transformer as input.
Moments of the posterior distribution. A consequence of knowing the sufficient statistic is finding the true posterior $p(\theta|x_{1:N})$, so we decode the moments of this posterior from the transformer embedding. The moments are functions of the mean and variance of the stream of data seen so far by the autoregressive predictor. In these conjugate models, it might be unsurprising for a transformer to encode the mean of the sequence it processes, because the optimal strategy for its loss function is to always predict the mean of the sequence that it sees so far. However, encoding the variance would not be directly related to this strategy and would support the argument that it infers sufficient statistics.
Out-of-distribution simulations. The analytical nature of the Bayesian predictor means that it is robust to datasets generated far from the prior – it would simply update its posterior based on the data. Thus, we probe the transformer on out-of-distribution (OOD) datasets that are generated from a distribution of the same form but with distinct hyperparameters.
@@
For Gaussian-Gamma hyperparameters $\alpha_{0},\beta_{0},\mu_{0},\lambda_{0}$ we choose $\{5,1,1,1\}$. The OOD dataset is instead generated with hyperparameters $\{2,1,5,1\}$, so it is centered on a different mean (5 vs. 1) and has a higher spread compared to the original data. An analogous design was used for the other exchangeable conjugate models, with details in Appendix A.2.
(a) Gaussian-Gamma (linear probe).
(b) Gaussian-Gamma (two-layer probe).
@@
Results
Results for all three models are shown in Figure 2, where the probe decodes the sufficient statistics. The remaining figures focus on the Gaussian-Gamma case (see Appendix A.3 for the other models). Figure 3(a) shows that the probe decodes the moments of the posterior distribution. Because the higher distribution moments are more volatile in value and less directly related to estimating the parameters $\theta$, we examine whether existing discrepancies are caused by an overly simple probe. We perform a second set of experiments where the probe has a hidden layer with ReLU activations (Figure 3(b)), showing stronger alignment.
Analyzing whether the transformer memorizes tokens
@@ Results
The transformer embedding (of size 128) can easily memorize a set of tokens, so we tested whether memorization is leading to the successful recovery of the sufficient statistics. We started by taking the 10th token embedding and probing the first 10 tokens’ sufficient statistics in each sequence, with high performance shown in Figure 8 in Appendix A.3. Then, we probed the token values themselves to look for memorization. Figure 4 suggests that memorization is generally absent. The 10th token embedding recovers the 10th token perfectly, but cannot recover the other 9 tokens. For the other 9 tokens, a correlation exists between probe results and true token values, but some level of correlation is expected because even seeing a single token can reveal information about this sequence’s generating distribution. However, the noise suggests that memorization is less of a goal than finding sufficient statistics.
@@ 4.2.1 Generative model
For an HMM generated data with $M$ sequences with $N$ tokens each, we formulate the generative process as, for a sequence ${x}$,
@@
where $C,V,{\pi}$ are initialized and denote, respectively, the number of classes, vocabulary size, and a list of probabilities on the number of classes to initialize the first latent state. $\gamma,\delta$ are scalar hyperparamters that are also initialized and fixed, and they represent the evenness of the samples from the Dirichlet distributions. $A$, $B$, and ${z}$, as a result, represent the transition matrix, the emission matrix, and the latent states, respectively.
We implement a forward-backward algorithm (Rabiner, 1989) to compute the posterior $p({z}_{n+1}|{x}_{1:n})$, and explore whether this distribution can be decoded from the transformer embedding on ${x}_{1:n}$.
@@ 4.2.2 Probing experiments
Like our experiments for exchangeable sequences, we use probes to explore whether an autoregressive transformer represents a predictive sufficient statistic for non-exchangeable sequences.
Decoding latent measures other than the predictive sufficient statistic
To further examine the role of the autoregressive objective in encoding a predictive sufficient statistic, we find other possible targets and test on whether they are more difficult to decode. We implement the Viterbi algorithm, which finds the most likely sequence of latent states given the whole observed sequence, and we use Viterbi’s output as our target. This is related to but distinct from the predictive sufficient statistic, $p(z_{n+1}|x_{1:n})$.
Implementational details
@@ Results
Based on the autoregressive objective, we hypothesized that a transformer trained on HMM sequences should encode the posterior distribution on latent states $p({z}_{n+1}|{x}_{1:n})$. Table 1 suggests that this is the case, reflected in terms of both accuracy of decoding and error in reproducing this posterior distribution (as measured by squared loss). Furthermore, decoding performance is better for the posterior than for several related quantities. It is more difficult to decode other quantities using the same embedding, or even to decode any of these quantities using the embedding on one token ahead, ${x}_{n+1}$. Several other quantities exhibit high volatility across random seeds, since the transition matrix has a greater influence on results when the tokens conditioned in the target posterior and the token used in the embedding have a mismatch.
Table 1: Probing target quantities in HMM dataset using different transformer token embeddings..
$\delta=0.5$ $\delta=1$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n}$ $90.8\pm 5.5\%$ $0.011\pm 0.012$ $90.4\pm 5.9\%$ $0.011\pm 0.01$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n+1}$ $66.6\pm 20.9\%$ $0.072\pm 0.042$ $65.9\%\pm 21.6\%$ $0.058\pm 0.035$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n}$ $53.2\pm 11.2\%$ $0.356\pm 0.043$ $50.8\pm 11.3\%$ $0.278\pm 0.038$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n+1}$ $86.5\pm 4\%$ $0.066\pm 0.034$ $82.2\pm 5.8\%$ $0.067\pm 0.036$
$\hat{{z}}_{n+1}$ ${x}_{1:n}$ $59.8\%\pm 12.1\%$ / $61.4\pm 14.7\%$ /
$\hat{{z}}_{n+1}$ ${x}_{1:n+1}$ $80.8\%\pm 7\%$ / $77.5\pm 8.8\%$ /
Note: The target $p({z}_{n+1}|{x}_{1:n})$ is a simplex vector found by running the forward-backward algorithm, and the target $\hat{{z}}_{n+1}$ is a scalar standing for the most likely latent class found by Viterbi algorithm. Average and standard deviation across 10 random seeds are reported.
4.3 Discrete hypothesis spaces
4.3.1 Generative model
We simulate a dataset where the data consist of a sequence of two-dimensional points uniformly sampled from a rectangular region in 2D space (cf. Tenenbaum, 1998). In this case, the hypothesis space $\mathcal{H}$ is defined as the set of all rectangles whose corner points are pairs of integers in $\{0,1,2,...,7\}$. For each sequence ${x}$, the generative process is,
$h_{\text{rect}}$ $\sim\mathcal{H},$
$x_{i}$ $\sim\textsc{Uniform}(h_{\text{rect}})\;\forall i.$
The generative process uniformly samples a rectangle $h_{\text{rect}}$ from the set of all rectangles $\mathcal{H}$. Then, each token in a sequence is sampled uniformly from the region defined by $h_{\text{rect}}$.
(a) A hypothesis space with equal-width rectangles.
(b) A hypothesis space with unequal-width rectangles.
Figure 5: Two discrete hypothesis spaces $\mathcal{H}$ used in our experiments. Any continuous rectangle contained within the axes (e.g., the red or the orange rectangle) is a valid hypothesis $h\in\mathcal{H}$. This results in 784 distinct hypotheses in each hypothesis space. The data consist of a sequence of points sampled uniformly from the target rectangle.

### v2

@@ 3.3 Probing embeddings to recover predictive sufficient statistics
These cases specify information that should be encoded in neural network embeddings. This sets up the second element of our approach, which is building probes to check this hypothesis. We focus on transformers (Vaswani et al., 2017), since they are widely used as language models. We denote the decoding target for each sequence $x_{1:n}$ by a vector $t_{n}$. Given a trained transformer, we decode $t_{n}$ by training a second model (a probe) to predict the target $t_{n}$ from the embedding $\phi_{n}$ of the corresponding document $x_{1:n}$. The embedding is defined to be the last layer embedding, which is what researchers and practitioners typically use as a document representation. The probe $g$ maps from the sequence embedding to the target. To ensure that the relevant statistical information is contained in the LM, not in the probe, we keep the probe simple by defining it as a linear layer with softmax activations: $g(\phi_{n})=\text{Softmax}(\text{Linear}(\phi_{n}))$. In each case, the probe uses the last-layer, last-token embedding as its input, unless otherwise specified.
4 Empirical analysis
We have identified cases where a predictive sufficient statistic is expected to be encoded by an autoregressive model. In this section, we use probing analyses on transformers to empirically validate this hypothesis. Ablation studies on memorization, parsimonious properties of embeddings, and decoding alternative quantities further support this idea. We use the Adam optimizer (Kingma and Ba, 2014) in all experiments. Code is available at github.com/zhang-liyi/llm-embeddings.
4.1 Simple exchangeable models
@@ 4.1.1 Generative model
Bayesian conjugate models
We start with three data generating distributions: Gaussian-Gamma, Beta-Bernoulli, and Gamma-Exponential models.
For $M$ Gaussian-Gamma sequences $\{x^{(k)}\}_{1\leq k\leq M}$ with $N$ tokens each, each sequence is i.i.d. generated by a mean and precision parameter sampled from a prior. For sequence ${x}^{(k)}$, the generative process is,
@@
where $\alpha_{0},\beta_{0},\mu_{0},\lambda_{0}$ are fixed hyperparameters. The full generative processes for the other two models are given in Appendix A.2.
The Gaussian-Gamma model generates from a Gaussian distribution with two unknown parameters, mean $\mu$ and precision $\tau$. The predictive distribution for the next token $x_{n+1}$ is
@@
The optimal Bayesian agent uses the same prior distributions as the Gaussian and Gamma that generate $\mu$ and $\tau$. It will analytically infer the ground truth posterior $p(\mu,\tau|x_{1:n})$ for any stream of data it sees, and use this posterior for predicting next tokens. A Bayesian agent can also use other suitable priors and converge to the optimal posterior. To be consistent with other exchangeable conjugate models, we denote $\theta=(\mu,\tau)$ to indicate latent variables whose posterior distribution is predictive sufficient.
Discrete hypothesis spaces
Each sequence in this dataset consists of two-dimensional points uniformly sampled from a rectangular region in 2D space (cf. Tenenbaum, 1998). The hypothesis space $\mathcal{H}$ is the set of all rectangles whose corner points are pairs of integers in $\{0,1,2,...,7\}$. The generative process uniformly samples a rectangle $h_{\text{rect}}$ from the set of all rectangles $\mathcal{H}$. Then, each token in a sequence is sampled uniformly from the region defined by $h_{\text{rect}}$.
@@
Figure 2: Probe recovery of transformer-learned sufficient statistic (blue) and ground truth sufficient statistic (red) on the y-axis, across 1000 test datapoints on the x-axis. In the plot above, the datapoints are sorted based on their ground truth sufficient statistic. The first row shows parameters probed in the non-OOD case (from left to right: Gaussian mean $\mu$, Gaussian precision $\tau$, Bernoulli mean, and Exponential mean). The second row shows the corresponding information in the OOD case.
(a) A hypothesis space with equal-width rectangles (left), and unequal widths (right).
(b) Colored by mean along y-axis of the generating rectangle.
(c) Colored by spread along y-axis of the generating rectangle.
Figure 3: (a): Two discrete hypothesis spaces $\mathcal{H}$ used in experiments. Any continuous rectangle contained within the axes (e.g., the red or the orange rectangle) is a valid hypothesis $h\in\mathcal{H}$. The data consist of a sequence of points sampled uniformly from the target rectangle. (b) and (c): Two-dimensional representation of embeddings of all validation datapoints (the setup is unequal width and sample size =50=50). The two subfigures show the same embeddings, colored by properties of the true generating rectangles.
Table 1: Probing results on discrete hypothesis spaces. In general, the probe achieves strong performance in recovering a 784-length vector, and performance increases as task difficulty decreases. Average and standard error across 10 random seeds are reported.
Equal Width Unequal Width
20 $87.3\pm 0.4\%$ $0.173\pm 0.005$ $66\pm 0.5\%$ $0.29\pm 0.004$
50 $99.5\pm 0.1\%$ $0.008\pm 0.001$ $88.5\pm 0.3\%$ $0.159\pm 0.003$
@@ 4.1.2 Probing experiments
For each dataset, the out-of-distribution (OOD) hyperparameters are chosen such that OOD data are centered on a different mean and has a different spread. The exact choices are detailed in Appendix A.3, along with hyperparameter sweep in Appendix A.5.
In discrete hypothesis spaces, we use rectangles with seven unit blocks on each side, resulting in a size-784 hypothesis space, i.e., 784 possible rectangles from which sequences are drawn. The unit lengths are either 1, or 0.4 and 0.16 alternating (Figure 3(a)).
Experimental results
We hypothesize that a transformer trained on these datasets should come to represent the sufficient statistics of the corresponding distribution.
Sufficient statistic. Results for the three Bayesian conjugate models are shown in Figure 2, where the probe successfully decodes the sufficient statistics $\theta$. For hypothesis spaces, we probe the distribution $p(h|x_{1:n})$, which is a length-784 simplex vector. Results are shown in Table 1. In general, the true hypothesis can be found with high accuracy, even though the number of classes is high.
Figure 3 visualizes how embeddings represent each hypothesis $h$ through Principal Components Analysis (PCA) that reduces the 128-dim embeddings to 2D, and the points are colored by properties of each rectangle. Embeddings are clustered into different regions based on the geometry of the true generating rectangle (Figure 3(b) and 3(c)), where, in 3(b) for instance, the top yellow clusters correspond to the upper-most rectangles, and dark blue clusters correspond to the lower-most rectangles. Additionally, the embeddings encode the distance between corners of the generating rectangle (Figure 3(c)). We also find the same color pattern (but rotated 90 degrees) when the horizontal instead of vertical axis is used to color the embeddings.
@@
Moments of the posterior distribution. A consequence of knowing the sufficient statistic is finding the true posterior $p(\theta|x_{1:N})$, and we show that the moments of this posterior can be decoded from the transformer embedding (see Figure 8 and 9 in the Appendix). The moments are functions of the mean and variance of the stream of observed data. In these conjugate models, it might be unsurprising for a transformer to encode the mean of the sequence it processes, because the optimal strategy for its loss function is to always predict the mean of the sequence that it sees so far. However, encoding the variance would not be directly related to this strategy and would support the argument that it infers sufficient statistics.
Out-of-distribution simulations. The analytical nature of the Bayesian predictor means that it is robust to datasets generated far from the prior – it would simply update its posterior based on the data. Thus, we probe the transformer on out-of-distribution (OOD) datasets that are generated from a distribution of the same form but with distinct hyperparameters. Figure 2, as well as 8(a) in the Appendix, suggest that the recovery of sufficient statistic and moments of the posterior distribution is generalizable to OOD cases, just as a Bayes-optimal agent would generalize.
Memorizing the context or storing sufficient statistics of the context?
@@
A possible confound behind the probe is that, instead of having a probe that uses the transformer-embedded sufficient statistics, the transformer may memorize a set of tokens, and the probe infers quantities of interest from these memorized inputs. We begin by making sure that the sufficient statistics over the first ten tokens can be decoded from the probe on the tenth token in the Gaussian-Gamma case (Figure 11 in Appendix A.5). Then, we probed the token values themselves to look for memorization. Figure 4 suggests that memorization is generally absent. The 10th token embedding recovers the 10th token perfectly, but cannot recover the other 9 tokens. For the other 9 tokens, a correlation exists between probe results and true token values, but some correlation is expected because even a single token can reveal information about the generating distribution. However, the noise suggests that the model is finding sufficient statistics rather than memorizing.
Embeddings are parsimonious
Our analysis so far has focused on whether embeddings contain information that allows decoding sufficient statistics. However, our hypothesis is stronger than this: since the sufficient statistics are enough to encode the relevant information from the data, a model need only represent that information. To explore whether transformers construct such parsimonious representations, we fixed the models trained on different distributions above and used a multi-layer perceptron (MLP) to directly predict their last-layer, last-token embeddings using the sufficient statistics of the training sequences as input.
Results show that in most cases the sufficient statistics capture well over 50% of variance in the embeddings (Table 6 in Appendix A.5). The results show that sufficient statistics capture a substantial amount of variance in transformer embeddings, establishing that the relationship between embeddings and sufficient statistics runs in both directions: we can decode sufficient statistics from embeddings, and we can predict embeddings from sufficient statistics.
Table 2: Probing target quantities in HMM dataset using different transformer token embeddings. The target $p({z}_{n+1}|{x}_{1:n})$ is a simplex vector found by running the forward-backward algorithm, and the target $\hat{{z}}_{n+1}$ is a scalar standing for the most likely latent class found by Viterbi algorithm. Average and standard error across 10 random seeds are reported.
$\delta=0.5$ $\delta=1$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n}$ $90.8\pm 1.7\%$ $0.011\pm 0.004$ $90.4\pm 1.9\%$ $0.011\pm 0.003$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n+1}$ $86.5\pm 1.3\%$ $0.066\pm 0.011$ $82.2\pm 1.8\%$ $0.067\pm 0.011$
$p({z}_{n+1}|{x}_{1:n})$ ${x}_{1:n+1}$ $66.6\pm 6.6\%$ $0.072\pm 0.013$ $65.9\pm 6.8\%$ $0.058\pm 0.011$
$p({z}_{n+1}|{x}_{1:n+1})$ ${x}_{1:n}$ $53.2\pm 3.5\%$ $0.356\pm 0.014$ $50.8\pm 3.6\%$ $0.278\pm 0.012$
$\hat{{z}}_{n+1}$ ${x}_{1:n}$ $59.8\pm 3.8\%$ / $61.4\pm 4.7\%$ /
$\hat{{z}}_{n+1}$ ${x}_{1:n+1}$ $80.8\pm 2.2\%$ / $77.5\pm 2.8\%$ /
@@ 4.2.1 Generative model
For an HMM generating data with $M$ sequences with $N$ tokens each, we formulate the generative process as
@@
where $C,V,{\pi}$ are initialized and denote, respectively, the number of classes, vocabulary size, and a list of probabilities on the number of classes to initialize the first latent state. $\gamma,\delta$ are scalar hyperparameters that are also initialized and fixed, and they represent the evenness of the samples from the Dirichlet distributions. $A$, $B$, and z{z}, as a result, represent the transition matrix, the emission matrix, and the latent states, respectively.
We use the forward-backward algorithm (Rabiner, 1989) to compute the posterior $p({z}_{n+1}|{x}_{1:n})$, and explore whether this distribution can be decoded from the transformer embedding on ${x}_{1:n}$.
@@ 4.2.2 Probing experiments
Implementation details
@@ Results
Our theoretical treatment suggests that the transformer should encode the predictive sufficient statistic $p(z_{n+1}|x_{1:n})$ or $p(z_{n}|x_{1:n})$ . The latter is the target used in belief state inference with transformers (Shai et al., 2024). However, there are other natural decoding targets that could be used. A simple target is the one-hot vector corresponding to the most likely hidden state $\hat{z}_{n+1}$. This is not a predictive sufficient statistic and does not encode the full information about the sequence relevant to future prediction. Our analysis thus suggests that it will thus provide a poorer match to the information contained in the embedding.
Table 2 suggests that the transformer encodes the predictive sufficient statistics. Furthermore, performance is better for our hypothesized measure (first and second row) than for related quantities (all other rows). Additionally, it is more difficult to recover the measure used in belief state inference (second row), although it is intuitively easier to learn: the target probability is over the same latent variable, but the emission of this latent variable is observed, unlike in our hypothesized measure where it is unobserved. This may be because if the embedding were to directly encode $p(z_{n}|x_{1:n})$, the transformer output layer after the embedding would additionally need to encode the transition matrix and additional integration computations, potentially making the task more challenging than encoding $p(z_{n+1}|x_{1:n})$.
4.3 Topic Models
We aim at recovering from language models the topic mixture $\theta_{i}$ that draws words in document $i$.
Table 3: Topic prediction by the autoregressive transformer (AT), Bert, LDA, and word-embedder (WE) on LDA-generated synthetic datasets, along with standard deviations across 3 random seeds. Hyperparameter $\alpha$ defines the dataset generation process, where a higher $\alpha$ means a more difficult task with underlying topics being more evenly distributed. AT and Bert have similar performance in the easiest setting, but AT performs well in harder settings where Bert performances worsen. End-to-end WE achieves stronger performance than language models, and LDA matches expectations by providing an upper bound in performance.
$\alpha$ Method Accuracy $\uparrow$ L2 loss $\downarrow$ Tot. var. loss $\downarrow$
AT $82.8\%\pm 0.5\%$ $0.041\pm 0.001$ $0.141\pm 0.001$
0.5 Bert $83.6\%\pm 1\%$ $0.036\pm 0.003$ $0.131\pm 0.005$
LDA $87\%\pm 0.6\%$ $0.029\pm 0$ $0.117\pm 0.001$
WE $85.8\%\pm 1.3\%$ $0.03\pm 0.001$ $0.119\pm 0.002$
AT $75.5\%\pm 0.8\%$ $0.044\pm 0.001$ $0.144\pm 0.001$
0.8 Bert $51.5\%\pm 1.7\%$ $0.111\pm 0.005$ $0.233\pm 0.011$
LDA $82.6\%\pm 0.5\%$ $0.036\pm 0.001$ $0.133\pm 0.004$
WE $80.9\%\pm 0.5\%$ $0.029\pm 0$ $0.116\pm 0.001$
AT $70.5\%\pm 1.6\%$ $0.045\pm 0.001$ $0.146\pm 0.003$
1 Bert $46.6\%\pm 3.3\%$ $0.1\pm 0.004$ $0.222\pm 0.006$
LDA $79.6\%\pm 1.4\%$ $0.045\pm 0.004$ $0.147\pm 0.006$
WE $79.4\%\pm 1\%$ $0.027\pm 0$ $0.113\pm 0.001$
(a) AT probe accuracies.
(b) Bert probe accuracies.
(c) Accuracy vs. neg. perplexity
Figure 5: Figure 5(a) and 5(b): control experiments showing AT (left) and Bert’s (middle) probe validation performance on synthetic data. For each AT and Bert, five models are trained and validated on five datasets, with each dataset generated by a distinct topic model. Colors show probe accuracy. A cell on row $i$ and column $j$ corresponds to model $i$ on dataset $j$, so the diagonal corresponds to a model on its own dataset. For AT, performance is only strong on the dataset with the same generating topic model, suggesting that the underlying statistical model, not the probe taking different word embeddings, is responsible for performance – a relationship that is also present for Bert but to a lesser degree. Figure 5(c): 20NG probe classification performance (accuracy) vs. negative perplexity measured at 100 different tokens. The dots are colored by the position percentile. Probe performance increases with lower perplexity.